# Strada-LLM: Graph LLM for traffic prediction

### v1

@@ III-B1 Hierarchical Feature Extractor
  • Global Graph Feature Extractor: While prompt-based networks are widely used in the existing methods [44], they lose the global structure information [14]. In order to keep the global graph structure information, we compute Laplacian embeddings [55]. The Laplacian operation leverages the eigenvectors of the complete graph and creates unique positional encodings [26] for each node. First, we define the graph Laplacian matrix  $L=D-A$ where $D=diag(d_{1},…,d_{N})$ is the degree matrix. We then compute the eigen decomposition of the normalized Laplacian as follows $L_{\text{norm}}=D^{-1/2}LD^{-1/2}=U\Lambda U^{T}$.
@@
  • Lag Features The tokenization strategy employed by Strada-LLM revolves around the creation of lagged features derived from historical traffic data. These features are constructed using predetermined sets of lag indices, such as quarterly, monthly, etc., also referred to as second-level frequencies, carefully chosen to capture various temporal patterns. Mathematically, this lagging operation can be expressed as a mapping $\boldsymbol{x}_{t}\longmapsto\boldsymbol{\lambda_{t}}\in\mathbb{R}^{|\mathcal{L}|\times F}$, where =$\mathcal{L}=\left\{1,...,L\right\}$ represents the set of lag indices. In this formulation, $\lambda_{t}[j]=x_{t-\mathcal{L}[j]}$, meaning that $\lambda_{t}[j]$ corresponds to the value of $x$ at $j$ time steps before $t$, as specified by the $j$-th element of $\mathcal{L}$. To put it another way, $\lambda_{t}[j]$ represents the $\mathcal{L}[j]$-th lagged feature when considering the time series $\boldsymbol{x}$ in reverse chronological order. This approach allows for a comprehensive encoding of temporal dependencies across various time scales. Therefore, to generate lag features for a specific context-length window $x_{1:C}$, where $C$ is the length of the context window, it is necessary to consider an expanded window that includes $L$ additional historical data points. Besides the lagged features, we incorporate date-time features covering all temporal features in our dataset, including second-of-minute, hour-of-day, and so forth, up to quarter-of-year, based on the time index $t$. It is important to note that the main purpose of these date-time features is to enrich the information set. In any time series, all but one of the date-time features will stay the same from one time step to another, allowing the model to intuitively understand the time series frequency. We use a total of $C$ date-time features, making the size of each token $|\mathcal{L}|+C$.
III-B2 LLM-based-Backbone
@@
Our backbone architecture is inspired by Mistral [38], a decoder-only transformer architecture. The extracted tokens by the sub-graph extractor are transformed by a shared linear projection layer that maps the features to the hidden dimension of the attention module. Inspired by [56], Strada-LLM includes pre-normalization by utilizing the RMSNorm [57] and Rotary Positional Encoding (RoPE) [58] for the representation of query and key blocks at each attention layer, similar to the approach described for Mistral [38].
@@
After passing through the causally masked transformer layers, the proposed method incorporates Flash-Attention-2 [59] to predict the parameters of the forecast distribution for the next timestamp, denoted as $\phi$. These parameters are generated by a parametric distribution head, as shown in Figure 2. The objective is to minimize the negative log-likelihood of the predicted distribution across all prediction times.
@@
At the inference stage, given a time series with the size of at least $L+C$, a feature vector can be formed and supplied to the model to determine the distribution of the subsequent timestamps. In this fashion, we can obtain many simulated future trajectories up to our chosen prediction horizon $T^{\prime}$ via greedy auto-regressive decoding [60]. In doing so, uncertainty intervals and confidence scores could also be valuable for downstream decision-making tasks.
@@
Figure 2: The proposed Strada-LLM architecture. The novelty lies in the K-hop sub-graph extractor module, which takes lag features in both the spatial and temporal dimensions. The transformer backbone is inspired by [56].
III-C Domain Adaptation
@@
In this section, we describe the techniques we adopt to adapt the Strada-LLM model to a new dataset, which might contain a graph of a new city. One of the important aspects of such a model is that it should be domain-agnostic. In the rapidly evolving landscape of deep learning, domain adaptation has emerged as a crucial technique for enhancing the generalizability of models across different but related domains. More specifically, domain adaptation focuses on the ability of a model trained in one domain (source) to adapt and perform well in another domain (target) with potentially different distributions. A prominent approach within this realm is low-rank adaptation [13], which aims to refine pre-trained models by adapting their parameters in a computationally efficient manner; indeed, a nearly linear operation will replace the $\mathcal{O}(n^{2})$ operation. This technique is particularly beneficial when working with LLMs, which, despite their impressive performance and versatility, pose significant challenges regarding computational resources and training time. In this paper, we adapt to the new distribution by tuning the distribution head, query, key, and value attention blocks.
Figure 3: Application of the LoRA method for domain adaptation. Frozen query, key, and value tensors will be fine-tuned by approximating low-rank matrices.
III-C1 Low-Rank Matrix Adaptation
@@
By leveraging domain adaptation strategies and Low-Rank Matrix Adaptation (LoRA) for fine-tuning,researchers and practitioners can effectively bridge the gap between domains, ensuring that powerful deep learning models maintain their efficacy across diverse applications [61]. Effectiveness of LoRA in the NLP domain has already been proven [13]. One of the purposes of this research is a demonstration of LoRA in the traffic forecasting problem. As illustrated in Figure 3, we apply LoRA for the fine-tuning process. The query, key, and value matrices will be fine-tuned in a low-rank approximation manner. Specifically, according to Figure 3, suppose the query, key, and value tensors are computed as follows:
$\small q=W_{q}h,k=W_{k}h,v=W_{v}h,$ (2)
@@
$\small\begin{split}\small q=W_{q}h+\Delta W_{q}h=W_{q}h+B_{q}A_{q}h\\ k=W_{k}h+\Delta W_{k}h=W_{k}h+B_{k}A_{k}h\\ v=W_{v}h+\Delta W_{v}h=W_{v}h+B_{v}A_{v}h,\end{split}$ (3)
where $h$ is the output of the previous layer, the low-ranked approximated update matrices $B_{q},B_{k},B_{v}$ are of dimensions $\mathbb{R}^{d\times r}$, $A_{q},A_{k},A_{v}$ are of dimensions $\mathbb{R}^{r\times k}$, and where the rank hyperparameter $r\ll min(d,k)$. By introducing these matrices, the online inference latency reduction is achieved [13]. For the initialization of the matrices $A$ we choose Gaussian initialization while the $B$ matrices are initialized to zero [13]. Following this approach enables us to perform domain adaptation for different datasets consisting of diverse distributions.
TABLE I: The comparison between PEFT and prompt-tuning approaches
# trainable params Memory Performance Flexibility to different architectures Prompt-Tuning [14] + LoRA [13] + + +
Table I presents a comparative analysis between Parameter-Efficient Fine-Tuning (PEFT) methods and prompt-tuning approaches, providing insights into their strengths and trade-offs. The table clearly highlights that prompt-tuning offers an advantage in terms of the number of trainable parameters, making it a lightweight alternative for model adaptation. However, it appears that parameter-efficient methods such as LoRA outperform prompt-tuning when considering memory efficiency, performance, and flexibility across different architectures. LoRA, in particular, demonstrates a clear edge in terms of both adaptability and overall resource management, suggesting its suitability for more complex and varied machine learning environments. These observations indicate that, while prompt-tuning may be appealing for its simplicity and parameter efficiency, methods like LoRA are likely more robust for scenarios requiring higher performance and architectural versatility.
III-D Loss Function
@@
$\small{\text{NLL}=-\sum_{x\in\mathcal{D}}log(P(y|x))},$ (4)
where $y$ is the true value, $x$ is the input data, $\mathcal{D}$ is the training set, and $P(y|x)$ is the predicted conditional probability of $y$ given $x$. In contrast to other loss functions, such as Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE) [14], which measure the difference between the predicted and true values directly, the negative log-likelihood measures the conditional probability of the true value under the model’s predictions. The network is fully optimized on the basis of the NLL loss during training. This probabilistic approach makes NLL particularly advantageous for models designed to capture the underlying probability distribution of the data. In the context of few-shot learning for a new geographical area, this characteristic enables the model to learn the intrinsic graph distribution effectively.
IV EXPERIMENTS
IV-A Dataset Details
@@
In this section, we assess the Strada-LLM model’s predictive capabilities using eight real-world datasets. The description of these datasets is presented in Table II. Most of these datasets pertain to traffic speed. For the sake of simplicity, traffic speed is utilized as the primary traffic metric in our experimental analysis; thus one can presume $F=1$.
TABLE II: Dataset Description
Dataset Data Type Number of Nodes Data Split Ratio(%)
Pretraining Datasets
PeMS03 [50] Traffic Speed 170 70, 20, 10
PeMS04 [50] Traffic Speed 307 70, 20, 10
PeMS07 [50] Traffic Speed 883 70, 20, 10
PeMS08 [50] Traffic Speed 170 70, 20, 10
Domain Adaptation
METR-LA [50] Traffic Speed 207 70, 20, 10
PEMS-Bay [50] Traffic Speed 325 70, 20, 10
PeMS07(M) [50] Traffic Speed 228 70, 20, 10
Brussels [62] Traffic Count 207 70, 20, 10
IV-B Training Strategy
IV-B1 Augmentation Techniques
We adopted the time series augmentation techniques $Freq\-Mix$[63] and $Freq\-Mask$ [63] to prevent over-fitting.
IV-B2 Hyperparameters
The hyperparameters of the Strada-LLM model mainly include: learning rate, batch size, training epochs, the number of layers, context length, embedding dimension per head, number of heads, rope scaling, and number of parallel samples for greedy decoding. We adopt Adam as the optimizer [64]. In the experiment, we follow the weight decay $81e-$ 8 and set the learning rate to 0.001 [64], the batch size to 32, and the training epochs to 150 for each dataset. We set $k=3$ in the sub-graph extractor. Furthermore, based on the fact that Strada-LLM is a probabilistic model, we take samples and compute medians to compare with point-wise models. The number of samples is set to $100$. In terms of complexity, our model contains 16 million parameters. The experiments are conducted on 2 Nvidia RTX 4090 GPUs.
IV-C Baselines

### v3

@@ 3.1.2. Global Graph Embedding
While prompt-based networks are widely used in the literature (Liang et al., 2024), they lose the global structure information (Li et al., 2024b). As demonstrated by  (Trillos et al., 2021), in order to keep the global graph structure information, we compute Laplacian embeddings. The Laplacian operation leverages the eigenvectors of the complete graph and creates unique positional encodings (Vaswani et al., 2023) for each node. First, we define the graph Laplacian matrix  $\textbf{L}=D-A$ where $D=\text{diag}(d_{1},\ldots,d_{N})$ is the degree matrix. We then compute the eigendecomposition of the normalized Laplacian as follows $L_{\text{norm}}=D^{-1/2}\textbf{L}D^{-1/2}=\textbf{U}\Lambda\textbf{U}^{T}$. Then we compute global embedding by non-linear mapping of U matrix. Reader can refer to Section 4.4.3 to discern the contribution of this representation extraction on the LLM forecasting ability. We denote $F^{\prime}$ as the embedding dimension.
3.1.3. Lag Extractor
Mathematically, this lagging operation can be expressed as a mapping $\boldsymbol{x}_{t}\longmapsto\boldsymbol{\lambda_{t}}\in\mathbb{R}^{|\mathcal{H}|\times F}$, where $\mathcal{H}=\left\{h_{1},h_{2},...,h_{H}\right\}$ represents the set of historical lag indices. In this formulation, $\lambda_{t}[j]=x_{t-\mathcal{H}[j]}$, meaning that $\lambda_{t}[j]$ corresponds to the value of $x$ at $h_{j}$ time steps before $t$, as specified by the $j$-th element of $\mathcal{H}$.
In our approach, we extend beyond simple lag feature extraction by employing a structured tokenization scheme to convert spatio-temporal slices of the road network graph into discrete tokens for the LLM-based backbone. Specifically, for each time step $t$, we first derive both global and local (k-hop) graph embeddings. These representations are then segmented within a sliding window, yielding embeddings that encapsulate the node-level or sub-graph-level features alongside their temporal contexts. Crucially, we add positional embeddings to each token so that the masked Transformer decoder can effectively attend to different time steps in a sequence-like manner. While lag features help capture short-term correlations, the tokenization scheme formalizes how each time window becomes a coherent input unit, thus ensuring the LLM’s attention mechanisms can leverage meaningful traffic patterns and dependencies across time and space.
@@ 3.1.3. Lag Extractor
Therefore, to generate lag features for a specific context-length window $x_{1:C}$, where $C$ is the length of the context window, it is necessary to consider an expanded window that includes $H$ additional historical data points. Besides the lagged features, we incorporate date-time features covering all temporal features in our dataset, including second-of-minute, hour-of-day, and so forth, up to quarter-of-year, based on the time index $t$. It is important to note that the main purpose of these date-time features is to enrich the information set. In any time series, all but one of the date-time features will stay the same from one-time step to another, allowing the model to intuitively understand the time series frequency. In the end, the HFE block’s output is a tensor in shape $\mathbb{R}^{C\times(1+M)\times|\mathcal{H}|\times(F+F^{\prime})}$.
3.2. LLM-based-Backbone
@@
Our backbone architecture is inspired by Mistral (Jiang et al., 2023), a decoder-only transformer architecture. The extracted tokens by the sub-graph extractor are transformed by a shared linear projection layer that maps the features to the hidden dimension of the attention module. Inspired by (Touvron et al., 2023), Strada-LLM includes pre-normalization by utilizing the RMSNorm (Zhou et al., 2021) and Rotary Positional Encoding (RoPE) (Su et al., 2023) for the representation of query and key blocks at each attention layer, similar to the approach described for Mistral (Jiang et al., 2023).
@@
After passing through the causally masked transformer layers, the proposed method incorporates Flash-Attention-2 (Dao, 2023) to predict the parameters of the forecast distribution for the next timestamp, denoted as $\phi$. These parameters are generated by a parametric distribution head, as shown in Figure 2. The objective is to minimize the negative log-likelihood of the predicted distribution across all prediction times.
@@
At the inference stage, given a time series with the size of at least $H+C$, a feature vector can be formed and supplied to the model to determine the distribution of the subsequent timestamps. In this fashion, we can obtain many simulated future trajectories up to our chosen prediction horizon $T^{\prime}$ via auto-regressive decoding (Bernardo et al., 2007). In doing so, uncertainty intervals and confidence scores could also be valuable for downstream decision-making tasks.
@@
Figure 2. The proposed Strada-LLM architecture. The core novelty lies in the K-hop sub-graph extractor module, which takes lag features in both the spatial and temporal dimensions.
3.3. Domain Adaptation
@@
In this section, we describe the techniques we adopt to adapt the Strada-LLM model to a new dataset, which might contain a graph of a new city. One of the important aspects of such a model is that it should be domain-agnostic. In the rapidly evolving landscape of deep learning, domain adaptation has emerged as a crucial technique for enhancing the generalizability of models across different but related domains. More specifically, domain adaptation focuses on the ability of a model trained in one domain (source) to adapt and perform well in another domain (target) with potentially different distributions. A prominent approach within this realm is low-rank adaptation (Hu et al., 2021), which aims to refine pre-trained models by adapting their parameters in a computationally efficient manner; indeed, a nearly linear operation will replace the $\mathcal{O}(n^{2})$ operation. This technique is particularly beneficial when working with LLMs, which, despite their impressive performance and versatility, pose significant challenges regarding computational resources and training time. In this paper, we adapt to the new distribution by tuning the distribution head, query, key, and value attention blocks.
(a) Strada-LLM (our method)
(b) UniST(Yuan et al., 2024)
Figure 3. Conceptual LLM adaptation approaches comparison
3.3.1. Low-Rank Matrix Adaptation
@@
By leveraging domain adaptation strategies and Low-Rank Matrix Adaptation (LoRA) for fine-tuning, researchers and practitioners can effectively bridge the gap between domains, ensuring that powerful deep learning models maintain their efficacy across diverse applications (Filatov and Kindulov, 2023). One of the purposes of this research is to propose a LoRA-based method for the temporal graph forecasting problem. As illustrated in Figure 3, we apply LoRA for the LLM fine-tuning process such as spatio-temporal patterns of the target data being learned. The query, key, and value matrices will be fine-tuned in a low-rank approximation manner. Specifically, according to Figure 3, suppose the query, key, and value tensors are computed as follows:
(2) $\small q=W_{q}h,k=W_{k}h,v=W_{v}h,$
@@
(3) $\small\begin{split}\small q=W_{q}h+\Delta W_{q}h=W_{q}h+B_{q}A_{q}h\\ k=W_{k}h+\Delta W_{k}h=W_{k}h+B_{k}A_{k}h\\ v=W_{v}h+\Delta W_{v}h=W_{v}h+B_{v}A_{v}h\end{split}$
where $h$ is the output of the previous layer, the low-ranked approximated update matrices $B_{q},B_{k},B_{v}$ are of dimensions $\mathbb{R}^{d\times r}$, $A_{q},A_{k},A_{v}$ are of dimensions $\mathbb{R}^{r\times k}$, and where the rank hyperparameter $r\ll min(d,k)$. By introducing these matrices, online inference latency reduction is achieved (Hu et al., 2021). For the initialization of the matrices $A$ we choose Gaussian initialization while the $B$ matrices are initialized to zero (Hu et al., 2021). Following this approach enables us to perform domain adaptation for different datasets consisting of diverse distributions. As shown in Figure 3(a), our adaptation approach is conceptually different than prompt-based models like UniST (Yuan et al., 2024). Based on the distribution-driven alignment loss function in the adaptation phase, global graph embedding, query, key, value projections, and distribution head are targeted to be tuned in the low-rank fashion. As reader can see in the experiment section 4.4.4, tuning the mentioned blocks is more lightweight than tuning prompt-based models like Figure 3(b). One might critique not applying low-rank tuning for other modules like Global Graph Embedding in Figure 3(a), which is addressed in Appendix  A.4.
Recent theoretical work shows that domain alignment can be bounded using the optimal transport distance functions (Yao et al., 2020). For instance, unsupervised distance adaptation approaches (Ramesh et al., 2024) have been adopted in modern LLMs. As shown in Figure  3(a), In Strada-LLM, the distribution head will be tuned. So, as another novel pillar of our model, during the adaptation phase, we minimize Kullback–Leibler divergence $D_{KL}(.)$ distance between multivariate distribution heads. As a matter of tractability, given a bounded $k$-variate $t$-distribution, the distance function can be derived as:
$D_{KL}(f_{d}(\cdot|\mu,\Sigma,\nu)||f_{d}(\cdot|\mu,\Sigma,\nu^{\prime}))=$
$=D_{KL}(f_{d}(\cdot|0,\mathbf{I},\nu)||f_{d}(\cdot|0,\mathbf{I},\nu^{\prime}))$
$=\int_{\mathbb{R}^{n}}f_{d}(\mathbf{x}|0,\mathbf{I},\nu)\log\frac{f_{d}(\mathbf{x}|0,\mathbf{I},\nu)}{f_{d}(\mathbf{x}|0,\mathbf{I},\nu^{\prime})}d\mathbf{x}$
Inspired by  (Villa and Rubio, 2018), this Kullback Leibler divergence can be written as:
(4) $D_{KL}(f_{d}(\cdot|\mu,\Sigma,\nu)||f_{d}(\cdot|\mu,\Sigma,\nu^{\prime}))=$
$\log\frac{K(d,\nu)}{K(d,\nu^{\prime})}-\frac{\nu+d}{2}\left[\Psi\left(\frac{\nu+d}{2}\right)-\Psi\left(\frac{\nu}{2}\right)\right]+$
$\frac{\nu^{\prime}+d}{2}K(d,\nu)\frac{\pi^{\frac{d}{2}}}{\Gamma(\frac{d}{2})}\int_{0}^{\infty}\left(1+\frac{t}{\nu}\right)^{-\frac{\nu+d}{2}}t^{\frac{d}{2}-1}\log\left(1+\frac{t}{\nu^{\prime}}\right)dt$
$\coloneqq{\mathcal{L}_{alignment}}$
where, $\small{K(d,\nu)=\frac{\Gamma(\frac{\nu+d}{2})}{\Gamma(\frac{\nu}{2})\sqrt{(\pi\nu)^{d}}}}$. The proof unfolds in Appendix A.2.
3.4. Loss Function
@@
(5) $\small{\text{NLL}=-\sum_{x\in\mathcal{D}}log(P(y|x))},\mathcal{L}_{total}=\lambda_{1}NLL+(1-\lambda_{1})\mathcal{L}_{alignment}$
where $y$ is the true value, $x$ is the input data, $\mathcal{D}$ is the training set, $P(y|x)$ is the predicted conditional probability of $y$ given $x$, and $\lambda_{1}$ is the loss component hyperparameter. In multivariate modeling, $NLL$ can be modeled differently. In Strada-LLM, we estimate the joint distribution, but further experiments are discussed in the ablation study 4.4.5.
4. EXPERIMENTS
Figure 4. Comparison of unified benchmark for different LLMs.
4.1. Datasets Details
@@
In this section, we assess the Strada-LLM model’s predictive capabilities using eight real-world datasets. The description of these datasets is presented in Table 1. Most of these datasets pertain to traffic speed. For the sake of simplicity, traffic speed is utilized as the primary traffic metric in our experimental analysis; thus one can presume $F=1$. Detailed training strategies, e.g., hyperparameters and experimental setup, are explained in Appendix A.5.
Table 1. Dataset Description
Dataset Data Type Number of Nodes Region Time Steps Pretraining Datasets PeMS03 (Fang et al., 2021) Traffic Speed 170 California 26208 PeMS04 (Fang et al., 2021) Traffic Speed 307 California 16992 PeMS07 (Fang et al., 2021) Traffic Speed 883 California 26208 PeMS08 (Fang et al., 2021) Traffic Flow 170 California 17856 Few-Shot METR-LA (Fang et al., 2021) Traffic Speed 207 California 34272 PEMS-Bay (Fang et al., 2021) Traffic Speed 325 California 52116 PeMS07(M) (Fang et al., 2021) Traffic Speed 228 California 12672 Brussels (Moghadas et al., 2024) Traffic Count 207 Brussels 12672 Zero-Shot Crowd (Yuan et al., 2024) Pedestrian Count $16\times 20$ Nanjing 9420 PeMS07(M) (Fang et al., 2021) Traffic Count 228 California 12672 Brussels (Moghadas et al., 2024) Traffic Count 207 Brussels 12672
4.2. Baselines
We compare the performance of the Strada-LLM model with spatio-temporal methods, GMAN  (Zheng et al., 2020), MTGNN  (Wu et al., 2020), GTS  (Shang et al., 2021), STEP  (Shao et al., 2022), and LLM-based methods which either are point-wise forecasters such as FlashST (Li et al., 2024b), UniST (Yuan et al., 2024) or probabilistic forecasters such as Lag-llama(Rasul et al., 2024). The details of the baseline methods are explained in Appendix A.1. To compare with the baselines, we adopt the MAE, RMSE, and MAPE prediction error metrics (Li et al., 2024b). Due to Strada-LLM being a probabilistic model, we also compare with Lag-Llama (Rasul et al., 2024) using the Continuous Ranked Probability Score (CRPS) metric, defined as:
$\small{CRPS(F,y)=\int_{-\infty}^{\infty}(F(x)-H(x-y))^{2}dx,}$
where $F(x)$ is the cumulative distribution function (CDF) of the forecasted distribution, $y$ is the observed value, and H(xy)H(x-y) is the Heaviside step function, which is 0 when $x<y$ and 1 when $x\geq y$.
4.3. LLM Forecasting Results
4.3.1. LLM Short-Term Forecasting
The short-term traffic prediction accuracy was evaluated across four datasets: PeMS-Bay, METR-LA, PEMS07(M), and Brussels. The last dataset is private. Strada-LLM, as an LLM should be benchmarked jointly for all datasets.
The prediction results for the datasets METR-LA and PEMS-Bay are reported in Tables 23, respectively. The ability to perform forecasting and being robust to new domains are important qualities for an LLM. As shown in Tables 23, the proposed LLM has competitive performance compared to STEP (Shao et al., 2022), which is the corresponding transformer-based fully-supervised method. Readers can refer to Appendix 10 about robustness analysis. On the other hand, compared to the best prompt-tuning method FlashST (Li et al., 2024b), shows that taking the global structure of the graph into account brings forth a definite advantage. Moreover, although Strada-LLM and STEP (Shao et al., 2022) are both transformer-based models, the introduction of normalization blocks like RMSNorm (Zhou et al., 2021) and Rotary Positional Encoding (RoPE) differentiates them. Additionally, while STEP (Shao et al., 2022) is an encoder-decoder model, Strada-LLM is a decoder-only model, which makes it optimized in terms of performance for inference (Fu et al., 2023). As we can see in Tables 23, the proposed Strada-LLM manages to outperform the baseline models on the aforementioned datasets in the longer prediction horizon more significantly. Specifically, on the PeMS-Bay dataset, our proposed model achieves an RMSE of 3.94 for 1-hour prediction, corresponding to an improvement of approximately $16\%$ compared to the prompt-tuning baseline (Li et al., 2024b).
Table 2. Traffic short-term forecasting comparison with the state-of-the-art methods on the METR-LA dataset
Baselines 15 min 30 min 60 min
[…]
Supervised
GMAN (Zheng et al., 2020) 2.80 5.55 7.41 3.12 6.49 8.73 3.44 7.35 10.07
MTGNN (Wu et al., 2020) 2.69 5.18 6.88 3.05 6.17 8.19 3.49 7.23 9.87
GTS (Shang et al., 2021) 2.67 5.27 7.21 3.04 6.25 8.41 3.46 7.31 9.98
STEP(Shao et al., 2022) 2.61 4.98 6.6 2.99 5.97 7.96 3.37 6.99 9.61
LLM
FlashST (Li et al., 2024b) 2.84 5.57 7.45 3.19 6.43 9.04 3.60 7.44 10.68
UniST (Yuan et al., 2024) 2.89 5.63 7.49 3.26 6.51 9.09 3.65 7.52 10.79
Strada-LLM(ours) 2.45 4.71 6.42 2.97 5.84 7.87 3.35 6.73 9.47
Table 3. Traffic short-term predictions comparison with the state-of-the-art methods on PEMS-Bay dataset
Baselines 15 min 30 min 60 min
[…]
Supervised
GMAN (Zheng et al., 2020) 1.34 2.91 2.86 1.63 3.76 3.68 1.86 4.32 4.37
MTGNN (Wu et al., 2020) 1.32 2.79 2.77 1.65 3.74 3.69 1.94 4.49 4.53
GTS (Shang et al., 2021) 1.34 2.83 2.82 1.66 3.78 3.77 1.95 4.43 4.58
STEP(Shao et al., 2022) 1.26 2.73 2.59 1.55 3.58 3.43 1.79 4.20 4.18
LLM
FlashST (Li et al., 2024b) 1.37 2.96 2.88 1.72 3.98 3.89 2.02 4.69 4.79
UniST (Yuan et al., 2024) 1.42 3.01 2.91 1.78 3.99 3.91 2.07 4.72 4.81
Strada-LLM (Ours) 1.28 2.77 2.61 1.53 3.14 3.37 1.76 3.94 4.09
Table 4. Long-Term prediction experiment on the PEMS07 (M) and Brussels dataset ($p\text{-value}<0.05$)
[…]
4.3.2. LLM Long-Term Forecasting
Followed by the setup of  (Yuan et al., 2024), we experiment with long-term forecasting for horizons 90, 120, and 150 minutes. The long-term traffic prediction accuracy was evaluated across two datasets: Crowd and PEMS07(M). We report the performance in Table 4. Consistently, Strada-LLM outperforms the LLM-based baselines. The successful performance can be attributed to the Graphical modeling of spatial covariates. Further visualizations are presented in the Appendix A.3.
Table 5. Few-shot learning experiment on the PEMS07(M) and Brussels dataset($p\text{-value}<0.05$)
[…]
4.3.3. LLM Zero-Shot Forecasting
The zero-shot traffic prediction accuracy was evaluated across three datasets: Crowd, Pems07(M), and Brussels. Table  7 presents the results of zero-shot capability of Strada-LLM compared to the other LLMs. Further visualizations are presented in the Appendix A.3. Strada-LLM is able to capture the spatio-temporal dependencies between connected nodes and prediction correlated trends, which is neglected by UniST (Yuan et al., 2024). To demonstrate the generalizability to the other urban datasets, Strada-LLM’s performance on the Crowd dataset (Yuan et al., 2024) is demonstrated in Figure 5, which outperforms other LLM baselines.
15min30min60min6.56.5777.57.5886.536.536.876.877.177.176.866.867.057.057.317.317.027.027.127.127.637.63HorizonsMAEOursUniST(Yuan et al., 2024)FlashST(Li et al., 2024b)
Figure 5. Zero-shot performance comparison on Crowd dataset.
4.3.4. LLM Probabilistic Forecasting
Table 6 reports the CRPS metric, likewise Lag-llama(Rasul et al., 2024), it has been implemented as the mean quantile loss. We observe that Strada-LLM consistently achieve strong few-shot and few-shot performance, obtaining either the best or second-best results for the Pems07 and Brussels datasets.
Table 6. Probabilistic forecasting experiment on the PEMS07(M) and Brussels dataset
PEMS07(M) Brussels
Baselines $\downarrow CRPS\downarrow$
15 min 30 min 60 min 15 min 30 min 60 min
Lag-llama (Rasul et al., 2024) 3.95 4.13 4.92 5.07 5.77 5.91
Strada-LLM(ours) 3.13 3.67 4.03 4.93 5.10 5.31
Table 7. Zero-shot learning experiment on the PEMS07(M) and Brussels dataset($p\text{-value}<0.05$)
[…]
4.4. Ablation Study
Table 8. Domain adaptation comparison for demonstrating selection of effective source dataset
Source $\rightarrow$ Target $Top^{k}$ Fine-tuning Ours Fully-Finetune Fully-Supervised
$downarrow MAE\downarrow RMSE\downarrow$ $downarrow MAE\downarrow RMSE\downarrow$
Metr-LA Pems-BAY 3.31 5.76 2.64 4.83 2.01 3.89 1.54 3.53
PeMS04 Pems-BAY 2.73 5.02 1.96 4.15 1.67 3.67 1.54 3.53
Table 9. The computation cost on the PEMS07(M) with the batchsize=64.
Method Parameters(M)
Trainable
Parameters(M)
Trainable
[…]
FlashST(Li et al., 2024b) 64.02 1.01 1.58
UniST(Yuan et al., 2024) 30.00 0.71 2.39
Strada-LLM (ours) 40.00 0.52 1.32
4.4.1. Domain Adaptation
We examine the adaptation capabilities of Strada-LLM by evaluating the gap when different datasets are used as source and target domains. Specifically, we trained Strada-LLM on one dataset (e.g., PeMS-04) and adapted to another (e.g., METR-LA) to assess its generalization across different types of traffic patterns. Table 8 presents the results for the domain adaptation process when pre-trained on a source dataset and adapted to the corresponding target dataset. The results presented in Table 8 indicate that there is a performance drop when the model is applied to a different target domain. This table can also address selecting efficient pre-training datasets to gain transferability advance for the targeted LLM. Take Pems-Bay as an example, selecting Pems04 as the pertaining dataset for LLM ends up the 14.07% superior performance in the adaptation phase compared to Metr-LA, because of the inherent closer distribution. Detailed ablated studies about the performance of LoRA in terms of the rank parameter are discussed in Appendix A.6.
Figure 6. Latent space embedding of the last layer (before the distribution head) visualized in a 2D space.
4.4.2. Model Interpretation
To better understand Strada-LLM, we visualize the latent space of the LLM representation for different datasets. By doing so, the disparity of dataset distributions and the reason for the difficulty of adapting to PeMS04 are revealed. The visualization is shown in Figure 6 which is obtained with T-SNE (Cai and Ma, 2022) in 2D dimensions. Our variational posterior distribution $P(\theta\mid X)$ can truly depict the underlying distribution.
4.4.3. Scalability of K-hop structure
Our experimental analysis explores Strada-LLM’s sensitivity to different neighborhood radii by examining performance across multiple k-hop values ($k\in\{1,3,5\}$). The results in Table 10 demonstrate a trade-off between prediction accuracy and neighborhood size, with larger radii yielding lower error rates. Comparative analysis presented in Table 6 shows that Strada-LLM achieves superior performance over Lag-Llama (Rasul et al., 2024), which did not take neighbors into account, primarily due to our model’s incorporation of spatial dependencies through the K-hop block structure, a feature absent in the baseline approach.
Table 10. Feature Scalability Analysis Benchmark
[…]
11 2.96 5.34 7.35
33 2.60 4.86 6.51
55 2.63 4.89 6.54
4.4.4. Computation Cost
We evaluate the performance of our proposed LoRA-based approach compared to existing LLMs. Results are demonstrated in Table 9. As can be seen, the lower error comes with the cost of involving more parameters. However, in contrast to UniST (Yuan et al., 2024), Strada-LLM’s adaptation outperforms the compression rate, from 1.58(%) to 1.32(%), which demonstrates its efficiency. For real-time applications, more lightweight models are being demanded, which remains the future direction. Performance over parameter analysis is illustrated in Appendix A.8.
4.4.5. Distribution modeling
We evaluate our model’s flexibility by deploying it across multiple LLM architectures. So, we report the performance in Table 11. Consistently, Strada-LLM outperforms the LLM-based baselines. The successful performance can be attributed to the Graphical modeling of spatial covariates.
Table 11. Ablation study on distribution modeling for PEMS07(M)
[…]
Univariate 2.81 5.63 6.81
Independent 2.69 5.37 6.73
Joint 2.60 4.86 6.51
4.4.6. LLM Cross-Compatibility
We evaluate our model’s flexibility by deploying it across multiple LLM architectures. Specifically, we utilized Mistral and Llama 2 as base models to assess the adaptability of our proposed method. The results are shown in Table 12.
Table 12. Ablation study on LLM backbone adaptability for PEMS07(M)
[…]
Llama2 (Rasul et al., 2024) 2.83 5.66 6.89
Mistral (Jiang et al., 2023) 2.60 4.86 6.51
5. Conclusion
This research proposes a probabilistic-based LLM for traffic forecasting called Strada-LLM, which injects the local graph structure implicitly into the input tokens. We utilize a subgraph extraction procedure to inject the graph structure into the transformer token inputs. On the one hand, we adopt a probabilistic transformer to predict the traffic data by sampling from the underlying distribution. On the other hand, we utilize a low-rank method for the transfer learning task. Consequently, we showcase its success in adapting to datasets with significant differences compared to the datasets used for pre-training. Finally, Strada-LLM is compliant to a variety of known LLM backbones. In summary, the Strada-LLM model successfully captures the spatial and temporal features from traffic data so that they can be applied to other spatio-temporal tasks. A potential avenue for future research could involve evaluating the expressiveness of the distribution head.
Appendix A Appendix
A.1. Baselines